If a/b =c/d=k ,3a3/3b-ac+c/+dc,find k in algebraic
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)


If a/b =c/d=k ,3a3/3b-ac+c/+dc,find k in algebraic
Problem not clear to me.

Answer by greenestamps(13198)



If you want help with your problem, it makes sense for you to take the time to present the problem clearly.

There is NO possible meaning of

"3a3/3b-ac+c/+dc"....

Try again